# Truly Assessing Fluid Intelligence of Large Language Models through Dynamic Reasoning Evaluation

### v1

@@ Abstract
11footnotetext: Equal contribution22footnotetext: Corresponding author.
@@ 1 Introduction
In recent years, large language models (LLMs) openai2024gpto1 ; DeepSeekAI2025DeepSeekR1IR ; 2024claude ; openai2024gpt4o ; yang2024qwen2 have achieved remarkable success across various applications, such as intelligent chatbots zhang2023chatgpt ; ouyang2022training and code generation chen2021evaluating ; nijkamp2023codegen . Models such as OpenAI’s O1 and DeepSeek’s R1 leverage substantial test-time computation to refine their reasoning processes, explore diverse strategies, and learn from previous errors, exhibiting a degree of cognitive flexibility and self-correction that closely mirrors aspects of human-like thinking. Such advanced reasoning abilities are critical for solving complex problems in disciplines hendrycks2021math ; cobbe2021gsm8k ; lewkowycz2022minerva and code competitions chen2021codex ; li2022alphacode , and are poised to revolutionize scientific discovery wu2024tamgen ; bran2024chemcrow . As such, there is an urgent need for a rigorous and principled evaluation framework to systematically track and quantify the reasoning intelligence of cutting-edge LLMs.
Figure 1: (a) Examples of the latent rule that models are expected to understand and apply. (b) Comparison with previous benchmarks, showing that our DRE-Bench aligns with cognitive levels and controls complexity via variables. (c) Leaderboard of LLM intelligence on DRE-Bench, with accuracy on the vertical axis and stability on the horizontal axis. Higher accuracy and lower variance indicate stronger fluid intelligence.
@@
Existing evaluation benchmarks in reasoning can be broadly categorized into two major types: crystallized intelligence cattell1963theory ; schipolowski2014nature and fluid intelligence cattell1963theory ; kent2017fluid . Crystallized intelligence refers to a model’s ability to apply accumulated knowledge to solve problems. Representative benchmarks such as AIME ye2025aime , GPQA rein2024gpqa , and SuperGPQA du2025supergpqa consist of high-difficulty discipline questions that require multi-step reasoning grounded in domain-specific knowledge. However, as models increasingly achieve expert-level performance on such knowledge-intensive tasks, it becomes increasingly important to evaluate their fluid intelligence, i.e., the ability to generalize beyond memorized content and reason in novel settings flanagan2007cattell . For example, benchmarks like ARC-AGI chollet2024arc and Raven’s Progressive Matrices raven2003raven assess this capacity through abstract reasoning tasks, i.e., infer the latent rule solely from provided input-output pairs and generalize it to predict correct outputs for novel inputs. Examples of the discovery of latent rules, such as frequency identification and category classification, are illustrated in Figure 1(a).
Although these benchmarks can measure fluid intelligence through task performance scores, they face several limitations. First, the latent rules are typically implicit and lack systematic alignment with human cognitive abilities. Most benchmarks chollet2019measure ; wu2025understanding ; yu2025stochastic present a collection of abstract reasoning tasks without explicitly aligning them with distinct dimensions or stages of human cognition primi2001complexity . As a result, it is difficult to determine how a model’s reasoning ability corresponds to specific levels of human-like intelligence. Second, each task is usually associated with a distinct latent rule, which introduces ambiguity in assessing the model’s true reasoning capability. Even if an LLM completes a task, it remains unclear whether it has genuinely understood and reasoned about the underlying rule or merely exploited superficial cues memorized during pre-training. Although recent efforts wu2025understanding ; yu2025stochastic have attempted to provide more fine-grained assessments of fluid intelligence, such as analyzing atomic operations wu2025understanding , these approaches still fall short in truly measuring fluid intelligence.
@@
To address these challenges, we propose a Dynamic Reasoning Evaluation benchmark, DRE-Bench, designed to assess the genuine fluid intelligence of large language models (LLMs). DRE-Bench is structured around a four-level cognitive hierarchy, from simple to complex reasoning. Each level assesses 3 latent rules specified by 3 abstract reasoning tasks. Each task includes multiple dynamic variants, all designed to evaluate the same underlying latent rule. In total, DRE-Bench provides over 3,000 abstract reasoning cases. This design enables fine-grained, cognition-aligned evaluation, and allows for a robust assessment of fluid intelligence by analyzing both accuracy and variance across rule-consistent task variants.
@@
Compared to existing benchmarks, DRE-Bench offers two key advantages as illustrated in Figure. 1(b). i) Cognition-aware task design. Unlike benchmarks such as ARC chollet2019measure and Atomic ARC wu2025understanding that present reasoning tasks without cognitive structure, DRE-Bench explicitly aligns each task with four human-like cognitive levels: attribute, spatial, sequential, and conceptual. This alignment provides good interpretability and allows mapping model behavior to specific cognitive capabilities. ii) Verifiable dynamic evaluation. To assess whether LLMs have truly grasped the underlying rules, DRE-Bench supports dynamic generation of multiple task instances by flexibly varying the latent rule-related variables. Due to the simple data format with numerical grids, all dynamic task samples and their answers can be generated deterministically via code. This ensures high reliability and correctness of evaluation data, a property that is difficult to guarantee in previous dynamic benchmarking approaches zhu2023dyval ; fan2023nphardeval ; zhu2024dynamic ; yang2024dynamic .
@@
We conduct comprehensive experiments on our proposed DRE-Bench with various LLMs, including general LLMs (models without thinking) such as GPT-4o openai2024gpt4o and Claude-3.7 2024claude , and reasoning LLMs (models with thinking) such as OpenAI-o1 openai2024gpto1 , DeepSeek-R1 DeepSeekAI2025DeepSeekR1IR , QwQ yang2024qwen2 , Skywork-OR1 skywork-or1-2025 , etc. The takeaways of our key findings are as follows:
@@
  • As the cognitive level of the reasoning tasks increases, model accuracy consistently declines, particularly for tasks involving physical concepts. Among them, OpenAI-o1 and DeepSeek-R1 demonstrate stronger performance and stability, while Claude 3.7 stands out in general LLMs as shown in Figure 1(c) and Section 4.2.
@@
Overall, the contributions of this paper are summarized as follows. 1) We propose an abstract reasoning benchmark with a cognition hierarchy, providing a more structural and comprehensive system to analyze the LLMs’ true fluid intelligence. 2) We develop a verifiable and scalable data engine to dynamically generate abstract reasoning data with various complexities, by designing a generator and solver for each task. 3) We perform comprehensive evaluations on a variety of popular LLMs, indicating that the existing LLMs still struggle to solve the reasoning problem of high cognitive levels. Existing LLMs may not have truly internalized the underlying reasoning rules, which highlights that they remain far from achieving true fluid intelligence. Our data and code will be available at https://github.com/yangyue5114/DRE-Bench.
@@ 2.1 Evaluation for Fluid Intelligence
There has been an abundance of attempts to define and measure the intelligence degree of existing large language models. Among them, the Abstraction and Reasoning Corpus(ARC)  chollet2019measure is regarded as a milestone, which defines that true intelligence should possess skill-acquisition efficiency. This concept attracted broad attention and led to many analytical studies wu2025understanding ; yu2025stochastic ; acquaviva2022communicating ; xu2023llms ; wang2023hypothesis ; wang2024speak .  wu2025understanding select some atomic abstract reasoning operations, and find that LLMs perform poorly on some atomic operations.  yu2025stochastic designed PHYSICO to evaluate whether LLMs really understand the physical phenomena they describe, by comparing language-format description and corresponding ARC format grid. However, existing abstraction reasoning benchmarks haven’t categorized tasks along cognitive dimensions, and can only provide a coarse-grained evaluation of LLMs’ reasoning ability. In addition, all these benchmarks are static, implying that they are highly susceptible to data contamination and only possess fixed complexity. Therefore, our work proposes DRE-Bench, a hierarchical cognitive dynamic benchmark on abstract reasoning. DRE-Bench can automatically generate data with varying levels of complexity, enabling comprehensive and fine-grained evaluation of LLM intelligence.
@@ 2.2 Dynamic Evaluation
Studies li2024latesteval ; li2024open ; yang2024dynamic have found that static benchmarks are highly prone to data contamination and have detected severe data contamination rates in some LLM benchmarks like wang2018glue ; wang2024mmlu . Moreover, their static nature implies a fixed level of complexity, making it difficult to adapt to evolving model capabilities. Therefore, some researchers have pioneered the exploration of dynamic evaluation on LLMs. Study (zhu2023dyval, ) proposed DyVal to dynamically generate test samples based on the graph structure to combat data contamination. Similarly, NPHardEval (fan2023nphardeval, ) generates new evaluation samples for NP-hard mathematical problems. To extend dynamic evaluation to more diverse NLP tasks, zhu2024dynamic further developed MPA, which employs LLM-based agents to transform existing problems into new ones. However, most of these dynamic evaluation methods are designed for general NLP tasks and are not applicable to more complex reasoning scenarios. More critically, the accuracy of their dynamically generated data is difficult to verify, leaving their reliability in constant doubt. In this work, we are the first to introduce a dynamic evaluation paradigm for abstract reasoning tasks. Our data generation process is code-verifiable, ensuring 100% reliability of the generated samples.
@@ 3.1 Constructing cognition-inspired abstract reasoning framework
Studies of fluid intelligence raven2003raven ; Carpenter2018 ; primi2001complexity indicate that the complexity of a reasoning problem may be related to the types of rules applied in the inductive reasoning process. The type of rules refers to different transformations applied to elements. Among them, the rule-type hierarchy proposed by Ricardo primi2001complexity represents a relatively comprehensive cognitive framework in psychology. This cognitive hierarchy categorizes inductive rule-type as four top-down levels, and for each level they outline several corresponding types of rule variables.
According to this cognitive hierarchy, we propose our abstract reasoning framework as shown in Figure 2. For the first-tier framework, we adopt four levels, namely (1) Attribute, (2) Spatial, (3) Sequential, and (4) Conceptual. Then, for each cognitive level, we summarize a series of related rule variables related to abstract reasoning tasks. Finally, for each rule variable, we design three sets of dynamic case generators to enable fine-grained evaluation of LLMs’ corresponding cognitive reasoning capabilities. The detailed dataset table is in Supplementary Materials.
Figure 2: Specific abstract reasoning tasks across our four cognitive levels. For each task example, we visualize two pairs of input and output, corresponding to two different values of the dynamic variable. The arrows are annotated with the value range of the variable, and darker colors indicate higher complexity.
Level-1: Attribute. In the attribute level, we follow the operational dimensions identified in cognitive psychology primi2001complexity , dynamically evaluating the reasoning capabilities of LLMs along three key rule types: size, count, and shape.
Level-2: Spatial. In the spatial level, drawing on psychological literature and ARC-related data, we designed a set of classic rules that comprehensively capture the notion of spatial reasoning, namely move, rotation, and symmetry. Specifically, for the “move” rule, we design dynamic data along five directional axes: up, down, left, right, and upper-right. For each direction, we generate five sets of data where the moving distance varies from 1 to 30. This enables a fine-grained assessment of the LLM’s understanding of both moving direction and distance. Similarly, for the “rotation” rule, we design two types of rotation axes, namely around an endpoint and around the center of objects. For each rotation setting, we generate dynamic variants by changing the rotation angle, ranging from 0 to 360 degrees. For the “symmetry” rule, we design tasks based on different axes of symmetry, covering three types: horizontal, vertical, and diagonal symmetry. For each type, the number of objects to be symmetrized can vary arbitrarily.
@@
Level-4: Conceptual. For Level-4, we focus on scientific concepts, which require not only high-level abstract reasoning but also the application of conceptual knowledge. Inspired by fundamental branches of physics yu2025stochastic , we construct three representative concepts: gravity, reflection, and expansion. To dynamically enhance the task complexity, we increase the times we apply physics rules.
@@ 3.2 Data generation framework
After determining the cognitive level, we proceed to select the specific rule to evaluate the LLM’s performance. To enable fine-grained and dynamic assessment of the model’s cognitive abilities, we design approximately three tasks for each rule. For example, the “move” rule includes five directional tasks such as up, down, left, right, and upper-right movement; the “rotation” rule includes three tasks: rotation around an endpoint, around the center, and rotation of objects. As shown in Figure 3, for each task, we identify its underlying constraint rule and construct a set of verifiable input-output pairs using tailored generators and solvers. This approach allows us to produce dynamic abstract reasoning cases in which variables vary within well-defined ranges.
Figure 3: DRE-Bench Data Generation Pipeline: (1) Annotators identify task-specific constraints and rules, writing a generator and solver function. (2) The generator produces dynamic input data based on different variable values. (3) The solver analyzes and computes ground-truth output for each input.
Constraint. First, for a given task case, our professional data annotators identify all case-relevant constraints as Figure 3(a), including <grid boundary>, <color assignment>, <specific structure>, <object placement>, and so on. Based on these constraints, the entire reasoning rule is encapsulated from scratch into a function pair—(generator, solver)—where dynamic variables are explicitly defined as function parameters. The generator produces diverse input grids according to different configurations of the dynamic variable parameter, while the solver derives the corresponding output grid for each input, ensuring 100% correctness of the generated data pair-(input grid, output grid).
Generator. In the second step, the generator is a task-specific function manually designed. As Figure 3(b), based on the case rules, it takes the corresponding dynamic variables as input parameters and generates the corresponding input grid. A random seed is embedded in the generator, enabling the scalable and reproducible generation of an unbounded number of diverse, constraint-satisfying samples. This automatically scalable data generation process can not only produce a large number of instances, but also ensure the trustworthiness of instances, achieving precise model evaluation.
Solver. Third, the solver is a paired function corresponding to the generator. Taking the input grid as one parameter, the solver encapsulates code that gradually parses the input grid and applies the corresponding rule to solve the task, ultimately producing the correct output grid. This step is executed entirely by code, ensuring 100% correctness of the output grid as ground-truth.

### v2

@@ 1 Introduction
Recently, large language models (LLMs)  (OpenAI, 2024b; DeepSeek-AI et al., 2025; Anthropic, 2024; OpenAI, 2024a; Yang et al., 2024a) have achieved remarkable success across various applications, such as disciplines (Cobbe et al., 2021; Lewkowycz et al., 2022), intelligent chatbots  (Zhang et al., 2023; Ouyang et al., 2022) and code generation  (Chen et al., 2021; Nijkamp et al., 2023). Models like OpenAI’s o1 (OpenAI, 2024b) leverage substantial test-time computation to refine their reasoning processes, learn from previous errors, and explore diverse strategies, exhibiting a degree of cognitive behavior that closely mirrors human-like thinking. As such, there is an urgent need for a principled evaluation framework to track and quantify the reasoning intelligence of cutting-edge LLMs systematically.
Figure 1: (a) Examples of the latent rule hidden in test cases. (b) Compared with previous benchmarks, our DRE-Bench demonstrates advantages in terms of hierarchy (cognition-aligned), scalability (code-generated), and dynamism (varying complexity). (c) Leaderboard of LLM intelligence on DRE-Bench, with accuracy on the y-axis and stability on the x-axis.
@@
Existing reasoning benchmarks can be broadly categorized into two major types: crystallized intelligence  (Cattell, 1963; Schipolowski et al., 2014) and fluid intelligence  (Cattell, 1963; Kent, 2017). Crystallized intelligence refers to models’ ability to apply accumulated knowledge to solve problems. Representative benchmarks such as AIME  (Ye et al., 2025), GPQA  (Rein et al., 2024), and SuperGPQA  (Du et al., 2025) which require multi-step reasoning grounded in domain-specific knowledge. However, as LLMs increasingly achieve expert-level performance on such knowledge-intensive tasks, the community gradually recognized that fluid intelligence—the ability to generalize beyond memorized content and reason in novel settings—is becoming increasingly important (Raven, 2003; Flanagan et al., 2007). In assessing the fluid intelligence of LLMs, ARC-AGI series (Chollet, 2019; Chollet et al., 2024) raise abstract reasoning tasks and is regarded as a milestone. Such tasks require LLMs to infer the latent rule solely from provided input-output training pairs and generalize it to predict correct outputs for novel testing inputs. Figure 1(a) illustrates two examples of such latent rules, frequency identification and category classification.
Although recent efforts (Chollet, 2019; Chollet et al., 2025) have attempted to measure the fluid intelligence of LLMs, such as analyzing atomic operations (Wu et al., 2025) and the stochastic parrot phenomenon (Yu et al., 2025), they face several limitations as shown in Figure 1(b). First, existing benchmarks usually comprise abstract reasoning cases whose latent rules are not linked with stages of human cognition  (Primi, 2001). Consequently, it is hard to tell what level of human-like intelligence a model has reached. Second, previous studies require manual annotation, which is labor-intensive and constrains benchmarks’ scalability and diversity of latent rules. Third, these benchmarks are inherently static, with each latent rule linked to only one or a few fixed cases. Such a static nature suffers from data contamination (Li et al., 2024a; Yang et al., 2024b), making it hard to determine whether the model truly understands the latent rule or merely memorizes it.
@@
To address these challenges, we propose a Dynamic Reasoning Evaluation benchmark, DRE-Bench, designed to assess the genuine fluid intelligence of large language models (LLMs). DRE-Bench is structured around a confirmed psychology hierarchy (Primi, 2001), with four cognitive levels ranging from simple to complex reasoning: Attribute, Spatial, Sequential, and Conceptual level. Each level contains 3 latent rules specified by several designed abstract reasoning tasks. Due to the simple data format of abstract reasoning tasks, we design a code-based generator and solver for each task, which can generate multiple dynamic variants with different levels of complexity. In total, DRE-Bench provides about 4K abstract reasoning cases. This framework enables a fine-grained, cognition-aligned evaluation of the abstract reasoning ability and allows for a robust assessment of fluid intelligence by analyzing both accuracy and variance across tasks with consistent latent rules.
@@
Compared to existing benchmarks, DRE-Bench offers three key advantages as illustrated in Figure 1(b). i) Cognition-aware task hierarchy. DRE-Bench presents reasoning tasks with a cognitive hierarchy, which explicitly aligns each task with four human-like cognitive levels. This alignment provides good interpretability and allows mapping model behavior to specific cognitive capabilities. ii) Human-Agent Collaboration Pipeline. For each latent rule, we employ LLM-driven agents to design code-based generators and solvers, which can produce input samples and corresponding answers accurately. To this end, our data generation pipeline achieves high correctness, efficiency, and scalability. iii) Dynamic evaluation. DRE-Bench supports dynamic generation of multiple task instances by flexibly varying the latent rule-related variables, obtaining extensive variants with different levels of complexity. This dynamic property helps avoid the data contamination issue that static datasets are prone to (Li et al., 2024a; b; Yang et al., 2024b). Therefore, we can precisely and comprehensively assess whether LLMs have truly grasped the underlying reasoning rules, further tracking the fluid intelligence of current LLMs.
@@
We conduct comprehensive experiments on DRE-Bench using a range of LLMs, including general-purpose models without explicit reasoning capabilities such as GPT-4o (OpenAI, 2024a) and Claude-3.7 (Anthropic, 2024), and reasoning LLMs (models with thinking) such as OpenAI-o1 (OpenAI, 2024b), DeepSeek-R1 (DeepSeek-AI et al., 2025), QwQ (Yang et al., 2024a), Skywork-OR1 (He et al., 2025), etc. The takeaways of our key findings are as follows:
@@
  • As the cognitive level of the reasoning tasks increases, model accuracy consistently declines, particularly for tasks involving physical concepts. Among them, OpenAI-o1 and DeepSeek-R1 demonstrate stronger performance and stability, while Claude 3.7 stands out in general LLMs. (Figure 1(c) and Section 4.2).
@@
Overall, the contributions of this paper are summarized as follows. 1) We propose an abstract reasoning benchmark with a cognition hierarchy, providing a more structural and comprehensive system to analyze the LLMs’ true fluid intelligence. 2) We develop a verifiable and scalable data engine to dynamically generate abstract reasoning data with various complexities, by designing a generator and solver for each task. 3) We perform comprehensive evaluations on a variety of popular LLMs, indicating that the existing LLMs still struggle to solve the reasoning problem of high cognitive levels. Existing LLMs may not have truly internalized the underlying reasoning rules, which highlights that they remain far from achieving true fluid intelligence.
@@ 2.1 Evaluation for Fluid Intelligence
There have been numerous attempts to define and measure the intelligence degree of existing large language models. Among them, the Abstraction and Reasoning Corpus(ARC)  (Chollet, 2019) is regarded as a milestone, which defines that true intelligence should possess skill-acquisition efficiency. This concept attracted broad attention and led to many analytical studies (Wu et al., 2025; Yu et al., 2025; Acquaviva et al., 2022; Xu et al., 2023; Wang et al., 2023; 2024a).  (Wu et al., 2025) select some atomic abstract reasoning operations, and find that LLMs perform poorly on some atomic operations.  (Yu et al., 2025) designed PHYSICO to evaluate whether LLMs really understand the physical phenomena they describe, by comparing language-format description and corresponding ARC format grid. However, existing abstraction reasoning benchmarks haven’t categorized tasks along cognitive dimensions, and can only provide a coarse-grained evaluation of LLMs’ reasoning ability. In addition, all these benchmarks are static, implying that they are highly susceptible to data contamination and only possess fixed complexity. Therefore, our work proposes DRE-Bench, a hierarchical cognitive dynamic benchmark on abstract reasoning. DRE-Bench can automatically generate data with varying levels of complexity, enabling comprehensive and fine-grained evaluation of LLM intelligence.
@@ 2.2 Dynamic Evaluation
Studies (Li et al., 2024a; b; Yang et al., 2024b) have found that static benchmarks are highly prone to data contamination and have detected severe data contamination rates in some LLM benchmarks like (Wang et al., 2018; 2024b). Moreover, their static nature implies a fixed level of complexity, making it difficult to adapt to evolving model capabilities. Therefore, some researchers have pioneered the exploration of dynamic evaluation on LLMs. Study (Zhu et al., 2023) proposed DyVal to dynamically generate test samples based on the graph structure to combat data contamination. Similarly, NPHardEval (Fan et al., 2023) generates new evaluation samples for NP-hard mathematical problems. To extend dynamic evaluation to more diverse NLP tasks, (Zhu et al., 2024) further developed MPA, which employs LLM-based agents to transform existing problems into new ones. However, most of these dynamic evaluation methods are designed for general NLP tasks and are not applicable to more complex reasoning scenarios. More critically, the accuracy of their dynamically generated data is difficult to verify, leaving their reliability in constant doubt. In this work, we are the first to introduce a dynamic evaluation paradigm for abstract reasoning tasks. Our data generation process is code-verifiable, ensuring 100% reliability of the generated samples.
@@ 3.1 Constructing cognition-inspired abstract reasoning framework
Studies about fluid intelligence (Raven, 2003; Carpenter et al., 2018; Primi, 2001) indicate that the complexity of a reasoning problem may be related to the types of rules applied in the inductive reasoning process. Among them, the rule-type hierarchy proposed by Ricardo (Primi, 2001) represents a relatively comprehensive cognitive framework in psychology. This framework categorizes inductive rule-type as four top-down levels, and proves the four levels form a true cognitive hierarchy: as from rule level 1 to 4, people impose qualitatively greater demands on abstraction, working memory, with reaction times and error rates also increasing. Therefore this categorization is suitable to assess the human-like fluid intelligence of LLMs.
According to this cognitive hierarchy of reasoning rule and corresponding rule variables, we propose our abstract reasoning framework as Figure 2. For the first-tier framework, we adopt four levels, namely (1) Attribute, (2) Spatial, (3) Sequential, and (4) Conceptual. Then, for each cognitive level, we summarize a series of related rule variables related to abstract reasoning tasks. Finally, for each rule variable, we design three sets of dynamic case generators to enable fine-grained evaluation of LLMs’ corresponding cognitive reasoning capabilities. The detailed dataset table is in Appendix C.
Figure 2: Specific abstract reasoning tasks across four cognitive levels. For each task, we visualize two pairs of input and output, corresponding to two different values of the dynamic variable. The arrows are labeled with variable ranges, with darker colors indicating higher complexity.
Level-1: Attribute. In the attribute level, we follow the operational dimensions identified in cognitive psychology (Primi, 2001), dynamically evaluating the reasoning capabilities of LLMs along three key rule types: size, count, and shape.
Level-2: Spatial. In the spatial level, drawing on psychological studies, we designed a set of classic rules that comprehensively capture the notion of spatial reasoning, namely move, rotation, and symmetry. Specifically, for the “move” rule, we design dynamic data along five directional axes: up, down, left, right, and upper-right. For each direction, we set the moving distance from 1 to 30. This enables a fine-grained assessment of the LLM’s understanding of both moving direction and distance. Similarly, for the “rotation” rule, we design two types of rotation axes, namely around an endpoint and around the center of objects. For each rotation setting, we change the rotation angle from 0 to 360 degrees. For the “symmetry” rule, we design tasks based on horizontal, vertical, and diagonal symmetry. For each type, the number of objects to be symmetrized can vary arbitrarily.
@@
Level-4: Conceptual. For Level-4, we focus on scientific concepts, which require not only high-level abstract reasoning but also the application of conceptual knowledge. Drawing inspiration from fundamental branches of physics (Yu et al., 2025), we introduce three representative concepts: gravity, reflection, and expansion. To further increase task complexity, we progressively intensify the application of these physical rules.
@@ 3.2 Data generation framework
After determining the cognitive level, we proceed to select the specific rule to evaluate the LLM’s reasoning performance. To enable fine-grained assessment, we design approximately three tasks for each rule. For example, the “move” rule includes five directional tasks: up, down, left, right, and upper-right movement. As shown in Figure 3, for each task, we identify its underlying constraint, then a code agent constructs a set of generators and solvers, upon human inspection, can be used to batch-produce input-output pairs. Such a human–agent collaboration pipeline can ensure scalability not only in the volume of data but also in the diversity of new rule.
Figure 3: DRE-Bench Data Generation Pipeline: (1) Professionals identify task-specific constraints and rules. (2) A CodeAgent collaborates with annotators to implement the generator and solver. (3) Different configurations are used to produce diverse cases.
Identifying Constraint. First, for a given task, professionals identify all case-relevant constraints, such as <grid boundary>, <color assignment>, <object placement>, and so on. These constraints, together with the corresponding rule, are then transformed into structured prompts, where a dynamic variable is explicitly defined. Each prompt subsequently invokes a code agent to generate two functions(i.e., a generator and a solver) parameterized by the dynamic variable.
Producing Generator and Solver. In the second step, an LLM-driven code agent is employed to implement the generator and solver functions for each task. Based on the rule and constraints encapsulated in the prompt (example in Appendix D), the code agent produces a generator that serves to generate the input grid, with a tunable parameter controlling the complexity of input cases. The paired solver is also implemented to parse the input grid and generate the corresponding ground-truth output grid. To ensure the correctness of the generator and solver, we predefine a set of parameter configurations to verify consistency between the input and output grids. If the generator–solver pair passes manual inspection, it is retained; otherwise, the code agent is re-invoked for refinement until a valid pair is produced. A random seed is embedded in the generator to enable scalable and reproducible generation of an unbounded number of diverse, constraint-satisfying samples.
Data Generation. Once the final generator and solver are established, for each rule, we can configure various parameters and different random seeds to generate batches of cases with varying levels of complexity. This data generation pipeline not only extends to large amounts of data with high correctness, but also ensures scalability to conveniently integrate new rules.